OpenStudy (anonymous):

i need help finding the y intercept on these two problem x+5y=25 and y=1/5x

OpenStudy (aroub):

first one: 5 second one : 0

OpenStudy (amistre64):

what defines a y intercept?

OpenStudy (angela210793):

5y=25-x y=5-x/5 y-intercept=5 and the second one is 0

OpenStudy (anonymous):

aroub! you are yellow!

OpenStudy (aroub):

lol i loved it!!! :)

OpenStudy (amistre64):

spose someone says the 1st is 3 and the 2nd is 12; how do you prove it?

OpenStudy (aroub):

ill prove him wrong...

OpenStudy (amistre64):

by what, restating that its 5 and 0?

OpenStudy (aroub):

oh i got what u mean! lol u want to me to write an explanation..!?

OpenStudy (amistre64):

a little something as a road map would be nice ;)

OpenStudy (aroub):

oh i will... next time!!! :D